Determine whether the following equations have a solution or not? Justify your answer.
m2 + m + 1 = 0

Answer by stanbon(75887)

Determine whether the following equations have a solution or not? Justify your answer.
m^2 + m + 1 = 0
determinant: = b^2 -4ac = 1^2-4*1*1 = -3
---------------
Since determinant is < 0, there are no Real Number solutions.
There are two unequal complex number solutions.
----------------
Cheers,
Stan H.